OpenStudy (anonymous):

Calculate the coordinates of the point on the curve y=(1-x)ˆ4 at which the gradient is -4.

OpenStudy (anonymous):

You will need to expand that equation, find dy/dx then set dy/dx equal to -4. Solving this equation for x should give you a set of x-coordinates that you can then plug back into your original equation to find the corresponding y-coordinates.

OpenStudy (anonymous):

I expanded the equation out but still couldn't get the x..I got a decimal

OpenStudy (jack1):

y = x^4 - 4x^3 + 6x^2 - 4x +1

OpenStudy (jack1):

you got this... yeah? ^^^^^^

OpenStudy (jack1):

and deriviative: dy/dx = 4x^2 - 12x^2 + 12x - 4

OpenStudy (anonymous):

I got -xˆ3+3xˆ2-3x

OpenStudy (jack1):

and as derivative is the formula of the gradient at any point in time so: when dy/dx = -4 -4 = 4x^2 - 12x^2 + 12x - 4

OpenStudy (jack1):

so it only works at x=0

OpenStudy (anonymous):

Isn't dy/dx just -4(1-x)ˆ3

OpenStudy (jack1):

so gradient = -4 at x = 0 you can sub x=0 into your initial eqn, or if ur really bright you can see tht the corresponding y value would equal c

OpenStudy (jack1):

"Isn't dy/dx just -4(1-x)ˆ3 " no dy/dx is actually 4 (x-1)^3

OpenStudy (jack1):

if u do it longhand and expand u cant make those mistakes easily though

OpenStudy (jack1):

ooo, shiny, cheers ;)

OpenStudy (anonymous):

No but it's a -x

OpenStudy (anonymous):

So 4x(-1x1)=-4

OpenStudy (jack1):

yeah, but you're not doing the chain or product rule correctly... but i cant really help with that, tht was @.Sam. and @campbell_st that were doing it that way

OpenStudy (jack1):

sorry, before derivative is : dy/dx = 4x^3 - 12x^2 + 12x - 4 i accidently put: dy/dx = 4x^2 - 12x^2 + 12x - 4

sam (.sam.):

Is it? \[ y=(1-x)^4\]

OpenStudy (jack1):

yep

sam (.sam.):

Chain rule? \[y'=4(1-x)^3 \cdot\frac{d}{dx}[1-x]\]

sam (.sam.):

You get \[y'=-4(1-x)^3\]

OpenStudy (jack1):

weird... i expanded and got something completely different

OpenStudy (jack1):

... u sure @.Sam. ...? i reduced mine down and got dy/dx is actually 4 (x-1)^3

sam (.sam.):

Mine is different, yours is 4(x-1)^3 mine is -4(1-x)^3

sam (.sam.):

Is just playing with the equations

OpenStudy (jack1):

ah... cool cheers

sam (.sam.):

cheers

OpenStudy (zzr0ck3r):

got to love them odd powers:)